find the length if a rectanglur lot with a perimeter of 122 meters if the length is 7 meters more than the width.
Answer by longjonsilver(2297)

Let width be x
so length is 7 more... ie it is x+7

So, perimeter = x+x+(x+7)+(x+7) and this equals 122
x+x+x+7+x+7 = 122
4x+14 = 122
4x = 108
x = 108/4
x = 27

so width = 27m
and length is therefore 34m

jon.